OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) and has a slope of 3 in point-slope form. x + 1 = 3(y – 5) x – 5 = 3(y + 1) y – 5 = 3(x + 1) y + 1 = 3(x – 5)

OpenStudy (anonymous):

i forget the equation for point-slope form

pooja195 (pooja195):

y-y1=m(x-x1) m=slope (x,y)

OpenStudy (abhisar):

\(\sf Y=mX +C\)

OpenStudy (sheraz12345):

y-y1=m(x-x1)

OpenStudy (anonymous):

so i have to use all of them?

pooja195 (pooja195):

@abhisat they want point slope

pooja195 (pooja195):

motown use wat i wrote

OpenStudy (anonymous):

i mean

OpenStudy (anonymous):

now that you wrote the equation its kind of self exploratory because y-(-1)=+1 which only one answer has

OpenStudy (lynfran):